Create the equivalent logic circuit of the following logic expression:

1. Q = (A + B) . (C +D)'

2. F1 = (A + BC') . D

3. Q = [(A + B)' . C] +B . C 


1
Expert's answer
2021-09-13T07:46:07-0400

1)Given logic expression is

"Q = (A + B) . (C +D)'\\=(A+B).(C'.D')"




2)Given logic expression is

"F_1 = (A + BC') . D"



3)Given logic expression is

"Q = [(A + B)' . C] +B . C\\=[A'.B'.C]+B.C"